need help!!

When five capacitors with equal capacitances are connected in series, the equivalent capacitance of the combination is 3.65mF. The capacitors are then reconnected sothat a parrallel combination of two capacitors is connected in series with a parallel combination of three capacitors. Calculate the equivalent capacitance of thiscombination. Enter your result in the unit of millifarad.
Answer #1
Capacitance of each capacitor c = cEquivalent capacitance when they are connected in series is C = 3.65 mFWe know C = c / 5From this c = 5C= 18.25 mFCombined capacitance of two capacitors when they are connected in series c ' = c/ 2=9.125 mFCombined capacitance of three capacitors when they are connected in parallel c " =c + c+ cc " = 3c= 54.75 mFWhen this combination is connectedseries withparallel combination of three capacitors then( 1/ C') = ( 1/c' ) + ( 1/ c" )From this equivalent capacitance C ' = ( c ' c " ) / ( c ' + c " )= 7.821 mF
answered by: Olili
Answer #2

When five capacitors with equal capacitances are connected in series, the equivalent capacitance of the combination is 7.03 mF. The capacitors are then reconnected so that a parallel combination of two capacitors is connected in series with a parallel combination of three capacitors. Calculate the equivalent capacitance of this combination. Enter your result in millifarads.

source: Openstax
answered by: Maya